OpenStudy (anonymous):

The vector u has a magnitude of 10. The x- and y-components of u are positive and equal and its z-component is 4. Determine x and y component of vector u .

OpenStudy (loser66):

\[\sqrt{x^2+y^2+z^2}=10\\x^2+y^2+z^2=100\\and~ x=y \]replace z =4 to solve for x, y.